Substitution
x+3y=7
3x=9-5y

Answer by friesr(113)

lets set first equal to y
-
3y=-x+7
-
y=-1/3x + 7/3
--
Sub in to second problem:
3x=9-5*(-1/3x+7/3)
-
x=-2
==
now put -2 in for x in first
-
-2+3y=7
-
3y=9
-
y=3
==
hope this helps